# Lumen: Unleashing Versatile Vision-Centric Capabilities of Large Multimodal Models

### v1

@@ 4.2 Results on Versatile Tasks
We evaluate our method on five vision-centric and vision-language tasks as shown in Table 1. We categorize existing approaches into three groups, namely “task-specific specialists”, “vision generalists” and “LMM generalists”, according to their functions and architectures. Task-specific specialists are customized models in different fields. They have diverse architectural designs and are limited to a single task. Vision generalists pursue handling multiple tasks with a unified architecture. To excel in fundamental visual perception tasks, they typically utilize a powerful vision encoder. LMM generalists aim to resolve every task in a conversational format. Focusing on improving the conversational quality, they adopt vision encoders proficient in multimodal content comprehension. For evaluation, we use COCO val set for both object detection and instance segmentation, COCO human body keypoints val set for pose estimation, and RefCOCOg val set for both visual grounding and referring segmentation. In the subsequent sections, we will analyze the experimental results based on the inherent characteristics of different tasks.
4.2.1 Object Detection & Instance Segmentation.
[…]
4.2.2 Pose Estimation.
Following the top-down paradigm employed by previous works [16, 10], we first crop a single object from the image using the bounding box. Therefore, the pose estimation is simplified to discovering keypoints of a single object. Since the LMM generalists do not perform this task, we only compare our method with vision generalists and task-specific specialists. As illustrated in Table 1, our method outperforms the vision generalist model Pix2Seq-v2 [10] with 0.6 AP. Meanwhile, our Lumen also achieves comparable performances with task-specific specialists.
4.2.3 Visual Grounding & Referring Segmentation.
[…]
@@
1 72.1 25.0
2 24.0 39.8
3 75.0 44.0
4 73.6 43.3
5 73.8 43.4
@@ 4.3 Ablation Studies
For efficiency, we train models for 10,000 iterations in ablation studies by default.
4.3.1 Multi-task Training.
@@
We use different task combinations for training the model, and the results are shown in Table 2. Based on the results, we have the following observations: (1) Compared with adopting datasets from all tasks (#5), discarding object detection data (#1) will reduce visual grounding performances. Similarly, removing visual grounding data (#2) also results in performance degradation on object detection. This demonstrates that object detection and visual grounding can benefit from each other. (2) Excluding pose estimation data leads to performance enhancements on both visual grounding and object detection (#3 vs #5). This might be caused by the data conflict problem addressed in many generalist models [55, 9]. (3) Excluding VQA data does not significantly affect model’s performances on both detection and visual grounding, which is reasonable as VQA data do not provide explicit object-level vision-language alignment cues.
4.3.2 V-L Dense Aligner Architectures.
@@
We ablate different architecture design choices as shown in Table 3(a), where “Conv.” indicates the operation of concatenating feature of [LOC] token with every image patch features and fusing them with two convolutional layers, and “Trans.” represents a light-weight transformer in our main method. It can be observed that substituting our transformer-based aligner with simple convolutional layers incurs significant performance degradation. This result indicates that complete vision-language interaction plays a pivotal role in dense prediction tasks like object detection, although this is less important in VQA tasks as demonstrated in LLaVA-1.5 [34].
4.3.3 Input Sizes.
@@
To examine the impact of various V-L dense aligner input resolutions, we resize the input image into three distinct sizes, concurrently adjusting the positional embeddings of the original CLIP ViT. As shown in Table 3(b), enlarging the input size from $256\times 256$ (default input size of CLIP ViT-L/14) to $448\times 448$ can achieve nontrivial performance boosts. However, further increasing the size to $896\times 896$ will harm the performances. This phenomenon indicates that excessively increasing the input size (to a scale dramatically different from its pretraining) will break the semantics of visual features, and thereby damaging the vision-language alignment.
4.3.4 Vision Encoders.
@@
Since the input size can not be seamlessly scaled up as discussed above, it is worth further exploring to leverage vision encoders naturally adaptive to high-resolution image inputs. Toward this end, we employ SAM ViT-H/16 [25], which takes $1024\times 1024$ images as inputs, as an additional vision backbone to provide high-resolution visual features for V-L dense aligner. However, as shown in Table 3(c), the utilization of SAM ViT-H/16 results in a significant performance deterioration. We posit that this phenomenon can be attributed to the inadequate alignment between the SAM visual encoder and the language modality. Building upon the aforementioned experimental findings and analysis, we think that embracing vision encoders capable of processing high-resolution image inputs without compromising semantic coherence can pave the way for further enhancing the vision-centric capabilities of LMMs.
Table 3: Ablation studies on model designs and hyper-parameters.
(a) Effect of different “V-L dense aligner” design choices.
[…]
Conv. 15.3 26.9 14.0
Trans. 27.4 43.4 27.1
(b) Effect of different input sizes for the dense alignment.
[…]
256$\times$256 25.0 38.9 24.8
448$\times$448 27.4 43.4 27.1
896$\times$896 22.3 35.2 21.8
(c) Effect of different pre-trained vision encoders.
@@
SAM ViT-H/16 12.3 20.0 12.6
CLIP ViT-L/14 27.4 43.4 27.1
(d) Effect of different $K$ value on dense prediction task.
@@
AP 25.7 26.4 27.4 26.6
[…]
(e) Effect of different numbers of training iterations.
@@
AP 27.7 29.7 31.8 33.0 33.9
[…]
4.3.5 Number of $K$ for Dense Prediction Tasks.
We also investigate the impact of selecting different values for $K$ on the dense prediction tasks, e.g., object detection. As shown in Table 3(d), when $K$ increases from 20 to 100, the model’s performance consistently improves, which can be attributed to the increased recall of true positives. However, as $K$ further increases from 100 to 200, the model’s performance degrades slightly due to the inclusion of false positives.
4.3.6 Training Epochs.
We record the variations in detection performance as the number of training iterations increases in Table 3(e). It can be observed that the model’s performance slowly increases as the training proceeds. Due to the limitation of computational resources, we did not train our model for a longer time. The convergence speed might be limited by the optimization difficulty of utilizing a single [LOC] token to query the entire image regions. A feasible solution to mitigate this problem is to employ more special tokens like [LOC] serving as object queries in DETR [5] with the Hungarian matching for label assignment, which can accelerate training and promote performances as proved in VisionLLM [55]. We do not adopt this design as it is unclear whether these tokens customized for object detection can generalize well to other tasks.
4.4 Generalization Evaluation
4.4.1 Generalization to Unseen Datasets.
To evaluate the generalization ability of our method, we perform zero-shot evaluation on PASCAL VOC2007 val set [14]. As illustrated in Table 4, our method demonstrates superior generalization ability than other generalist models. It is worth noting that compared with InstructCV which also inherits the enormous word vocabulary of LLM, our method outperforms it on VOC zero-shot evaluation by 17.0 AP. Besides, even compared with specialist models trained on VOC dataset (i.e, RetinaNet and Faster R-CNN in Table 4), our method still achieves comparable performances.
4.4.2 Generalization to Unseen Tasks.
To prove that the heatmap produced by our first-stage task-agnostic training is a powerful intermediate representation for scaling up to more vision-centric tasks, we apply simple decoding rules on the heatmap to make our Lumen support the object counting task [42]. Specifically, we first select 100 peak points following the operations stated in Sec 3.2. Then we filter the points with low confidence scores, and calculate the number of remaining points as the final prediction results. We illustrate user instructions, predicted heatmaps, selected points, as well as the ultimate predictions and ground-truth annotations as shown in Fig 4, where all images come from FSC-147 [42] dataset. It can be seen that our Lumen can make the correct prediction even without specifically training on the object counting task.
Table 4: Generalization ability evaluation on object detection task. $\dagger$ indicates that VOC is not included in our training data, and thereby used for cross-dataset generalization evaluation. Method COCO ${}^{\dagger}$ RetinaNet [31] - 77.3 Faster R-CNN [44] - 80.4 Pix2Seq-v2 [10] 57.4 38.5 InstructCV [15] 48.5 61.7 Lumen (Ours) 51.2 78.7
5 Conclusions
In this paper, we present Lumen, a Large multimodal model with versatile vision-centric capabilities enhancement. Within the Lumen, we employ a decoupled design to initially promote learning task-agnostic vision-language dense alignment, and subsequently collaborate with task-specific decoding operations to address diverse tasks. Thanks to this design, our Lumen not only significantly broadens the range of visual tasks that existing LMM generalists can tackle, but also exhibits remarkable generalization capabilities to unseen datasets and tasks.

### v2

@@ 4.2 Results on Versatile Tasks
We evaluate our method on vision-centric and vision-language tasks as shown in Table 1 and 2. We categorize existing approaches into three groups, namely “task-specific specialists”, “vision generalists” and “LMM generalists”, according to their functions and architectures. (1) Task-specific specialists are customized models in different fields. They have diverse architectural designs and are limited to a single task. (2) Vision generalists pursue handling multiple tasks with a unified architecture. To excel in fundamental visual perception tasks, they typically utilize a powerful vision encoder or additional designs (e.g, data augmentations or decoding strategies) adaptive to vision-centric tasks. (3) LMM generalists aim to resolve every task in a conversational format. Focusing on improving the conversational quality, they adopt vision encoders proficient in multimodal content comprehension. In line with these methods, our Lumen pursues unleashing the inherent vision-centric capabilities of LMMs.
Object Detection & Instance Segmentation. Object detection and instance segmentation require the model to make dense predictions across the image, therefore they pose great challenges in capturing fine-grained object cues. COCO val set is used for evaluation. (1) Compared with other LMM generalists, our Lumen surpasses them by clear margins, achieving a 15.6 AP50 boost over Griffon [12] and a 4.7 AP50 boost over InstructCV [61]. This indicates the effectiveness of our method in discovering dense object cues. (2) Compared with vision generalists and task-specific specialists, our method further approaches their performances. We analyze that the performance gap might originate from two major aspects. First, we use the input size of $448\times 448$, which is much smaller than these competitors, for example, $1024\times 1024$ in [13, 60]. Second, we do not introduce DETR-like object decoding techniques as done in [36, 60].
Pose Estimation. We employ COCO human 2D body keypoint val set for evaluation. Following the top-down paradigm employed by previous works [50, 13], we first crop a single object from the image using the bounding box. Therefore, the pose estimation is simplified to discovering keypoints of a single object. Since the LMM generalists do not perform this task, we only compare our method with vision generalists and task-specific specialists. As illustrated in Table 1, our method outperforms the vision generalist model Pix2Seq-v2 [13] with 2.4 AP. Meanwhile, our Lumen also achieves comparable performances with task-specific specialists.
Visual Grounding & Referring Segmentation. Compared to object detection and instance segmentation, visual grounding and referring segmentation underscore the language comprehension ability. Here, we report the results on RefCOCOg val set for comparison because the referring expressions in it are more diverse and abundant than those in RefCOCO and RefCOCO+. We also provide complete results on these three benchmarks in the Appendix B. As illustrated in Table 1, our method achieves better performances than Shikra [31] and Griffon [12] on visual grounding task, with AP50 83.6 vs 82.3 and 83.6 vs 81.5, respectively.
Visual Question Answering. To examine general visual understanding and instruction following of our model, we follow common practices and employ MMBench [62], MME [63], SEEDBench [64], MMMU [65] and MathVista [66] for evaluation. As shown in Table 2, our Lumen achieves VQA results comparable with state-of-the-art LMMs while extending versatile vision-centric capabilities.
Table 3: Ablation studies on model designs.
(a) Effect of different “V-L dense aligner” design choices.
Arch. AP AP50 AP75
Conv. 18.4 30.2 15.7
Trans. 28.4 45.0 28.1
(b) Effect of the pretrained mask decoder. cIoUs are reported.
Mask Dec. RC RC+ RCg
TransDec. 65.8 58.6 62.0
SAM 66.6 59.2 61.5
(c) Effect of LMM baseline.
Baseline AP AP50 AP75
LLaVA-v1.0 24.0 39.0 23.0
LLaVA-v1.0* 26.8 43.2 26.0
LLaVA-v1.5 28.4 45.0 28.1
Table 4: Effect of multi-task training. AP50 on RefCOCOg val set and COCO val set are reported.
@@
1 72.3 26.3
2 24.6 41.2
3 75.8 45.7
4 75.0 44.8
5 74.8 45.0
Table 5: Effect of the training recipe on VQA performances. MMBench is used for evaluation [^3] Please refer to the official benchmark for the specific meaning of each term..
# Phase 1 Phase 2 AR CP FP-C FP-S LR RR Overall
1 69.3 74.3 54.5 66.6 28.0 55.7 62.5
2 67.3 80.0 49.0 67.9 28.8 58.3 63.7
3 67.3 77.0 58.7 67.2 33.0 56.5 64.2
@@ 4.3 Ablation Studies
For efficiency, we default to training the model for 10,000 iterations in the first phase of task-agnostic learning. This protocol is standard for our ablation studies
@@
Multi-task Training. We use different task combinations for training the model, and the results are shown in Table 4. Based on the results, we have the following observations: (1) Compared with adopting datasets from all tasks (#5), discarding object detection data (#1) will reduce visual grounding performances. Similarly, removing visual grounding data (#2) also results in performance degradation on object detection. This demonstrates that object detection and visual grounding can benefit from each other. (2) Excluding pose estimation data leads to performance enhancements on both visual grounding and object detection (#3 vs #5). This might be caused by the data conflict problem addressed in many generalist models [36, 67]. (3) Excluding VQA data (i.e., discarding Phase 2) does not significantly affect model’s performances on both detection and visual grounding, which is reasonable as VQA data do not provide explicit object-level vision-language alignment cues.
Training Recipe. We ablate the effects of 2-phase training strategy to inspect the contribution of vision-centric data on VQA tasks. In Table 3, we list detailed scores on diverse aspects covered by MMBench. #1 is initialized from LLaVA’s weights, without any tuning. By comparing #2 and #3, we observe that FP-C (i.e, Cross-instance Perception) improves remarkably, which indicates that incorporating vision-centric data in Phase 1 can promote multi-instance perception in VQA (as well as the overall performance). Scores of Phase 1-only are not reported as the model trained without rich instruction data tends to respond template answers like “Sure, the location is [LOC].’’
@@
V-L Dense Aligner Architectures. We ablate different architecture design choices as shown in Table 3(a), where “Conv.” indicates the operation of concatenating feature of [LOC] token with every image patch features and fusing them with two convolutional layers, and “Trans.” represents a light-weight transformer in our main method. Substituting our transformer-based aligner with simple convolutional layers incurs significant performance degradation. This result indicates that complete vision-language interaction is crucial in promoting explicit dense vision-language alignment.
Pretrained Mask Decoder. We adopt the mask annotations from RefCOCO (RC), RefCOCO+ (RC+) and RefCOCOg (RCg) to train a lightweight transformer-like mask decoder from scratch (TransDec.) and report cIoU on their val set to ablate the effect of leveraging pretrained SAM decoder in Table 3(b). Although SAM outperforms TransDec., the narrow performance gap suggests that the primary challenge in advancing segmentation performances within our framework is the need to improve the dense vision-language alignment in the learned heatmap.
LMM Baseline. We ablate the effect of different LMM baselines on vision-centric task as shown in Table 3(c). Since LLaVA-v1.5 employs CLIP ViT-L/14-336 as a stronger vision backbone, we use it to substitute the original CLIP ViT-L/14-224 in LLaVA-v1.0, denoted as “LLaVA-v1.0*” in Table 3(c), to respectively ablate the effects of stronger vision backbone and multimodal language model. It can be proved that both stronger vision backbone and multimodal language model can benefit dense vision-language alignment.
Table 6: Generalization ability evaluation on object detection. $\dagger$ indicates that VOC is not included in our training data, and thereby used for cross-dataset generalization evaluation.
Method COCO VOC
RetinaNet [68] - 77.3
Faster R-CNN [51] - 80.4
Pix2Seq-v2 [13] 57.4 38.5
InstructCV [61] 48.5 61.7
Lumen (Ours) 53.2 77.9
4.4 Generalization Evaluation
Generalization to Unseen Datasets. To evaluate the generalization ability of our method, we perform zero-shot evaluation on PASCAL VOC2007 val set [69]. As illustrated in Table 6, our method demonstrates superior generalization ability than other generalist models. It is worth noting that compared with InstructCV which also inherits the enormous word vocabulary of LLM, our method outperforms it on VOC zero-shot evaluation by 16.2 AP. Besides, even compared with specialist models trained on VOC dataset (i.e, RetinaNet and Faster R-CNN in Table 6), our method still achieves comparable performances.
Generalization to Unseen Tasks. To prove that the heatmap produced by our first-stage task-agnostic training is a powerful intermediate representation for scaling up to more vision-centric tasks, we apply simple decoding rules on the heatmap to make our Lumen support the object counting task [70]. As shown in Fig 4 our Lumen can make the correct prediction even without specifically training on the object counting task.
5 Conclusions
In this paper, we present Lumen, a Large multimodal model with versatile vision-centric capabilities enhancement. With the task-agnostic and task-specific design, Lumen significantly broadens the range of visual tasks that existing LMM generalists can tackle, and also maintains general visual understanding and instruction following capabilities.
Appendix
Appendix A Implementation Details
A.1 Detailed Design of V-L Dense Aligner
Figure 5: Detailed designs of “V-L Dense Aligner”. [LOC] is the special token output from the LMM. [M] is the special token used for predicting the matching probabilities (i.e., heatmaps) in the first class-agnostic matching stage in the main paper. [H] and [W] are special tokens used as a simple box decoder for predicting additional height and width in the second class-specific decoding stage in the main paper. We use dotted lines to illustrate that these tokens are added in the second stage.
The detailed design of the transformer-based V-L dense aligner is shown in Fig 5. Motivated by SAM [34], we regard [LOC] token as a prompt token, which carries semantics conveyed in the instruction and is used to guide different output tokens, namely [M], [H] and [W], to generate corresponding predictions (i.e., heatmap, height and width in Fig 5). Specifically, the prompt token and output tokens are first interacted with a self-attention layer. Then, the resulting outputs are interacted with image embeddings through a sequence of cross-attention layers and a feed-forward network as shown in Fig 5 to perform dual attention. The above process is repeated twice. The final outputs are produced by performing an element-wise multiplication of the features of output tokens with the features of the image embeddings. For the sake of clarity and conciseness, we have omitted channel compression steps from Fig 5.
As for the output tokens, in the first stage, we only use [M] to generate the predicted heatmap. In the second stage, we add [H] and [W] to predict the height and width maps under the supervision of L1 loss as done in [40].
Table 7: Performance comparison with state-of-the-art specialists and generalists on the visual grounding task.
Method RefCOCO RefCOCO+ RefCOCOg
val test-A test-B val test-A test-B val test
Specialists
UNINEXT [71] 92.64 94.33 91.46 85.24 89.63 79.79 88.73 89.73
G-DINO-L [72] 90.56 93.19 88.24 82.75 88.95 75.92 86.13 87.02
Generalists
VisionLLM-H [36] 86.70 - - - - - - -
OFA-L [20] 79.96 83.67 76.39 68.29 76.00 61.75 67.57 67.58
Shikra-7B [31] 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
Lumen-7B (Ours) 88.59 92.06 83.55 80.35 87.59 72.20 83.62 84.44
Appendix B Complete Results on Visual Grounding and Referring Segmentation
We provide the complete performance comparison with existing approaches on visual grounding and referring segmentation tasks as shown in Table 7 and Table 8. We will analyze the results in the following parts.
B.1 Visual Grounding.
We report the visual grounding performances of our Lumen on RefCOCO, RefCOCO+ and RefCOCOg datasets in Table 7. First, compared with VisionLLM-H, a generalist that can also perform object detection and instance segmentation like us, our method outperforms it on the RefCOCO val set even without using a strong vision encoder, e.g., InternImage-H used in VisionLLM-H. Besides, compared with OFA-L and Shikra-7B, which reformulate the box prediction as a sequence generation manner, our method can achieve better or comparable performances with them. It indicates that our predicted heatmap can capture the correspondence between visual regions and complex language concepts.
B.2 Referring Segmentation.
We report the referring segmentation performances of our Lumen on RefCOCO, RefCOCO+ and RefCOCOg datasets in Table 8. First, compared with other generalists, our Lumen surpasses them across most benchmarks by clear margins. Besides, it is worth noting that we do not use any pixel-level supervision (i.e., segmentation masks) throughout our training process, but our method still achieves comparable performances with specialist models on the challenging RefCOCOg benchmark.
Table 8: Performance comparison with state-of-the-art specialists and generalists on the referring image segmentation task.
Method RefCOCO RefCOCO+ RefCOCOg
val test-A test-B val test-A test-B val test
Specialists
MCN [73] 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4
LAVT [58] 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
ReLA [59] 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0
LISA [33] 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5
Generalists
X-Decoder [74] - - - - - - 64.6 -
Unified-IO [22] 46.4 46.1 48.1 40.5 42.2 40.2 17.3 -
InstructDiff [50] 61.7 65.2 60.2 46.6 52.3 39.0 67.4 -
Lumen (Ours) 70.7 75.5 66.2 62.8 69.6 55.4 65.1 67.0
Appendix C More Ablation Studies
C.1 Input Sizes
@@
To examine the impact of various V-L dense aligner input resolutions, we resize the input image into three distinct sizes, concurrently adjusting the positional embeddings of the original CLIP ViT. As shown in Table 9(a), enlarging the input size from $336\times 336$ (default input size of CLIP ViT-L/14) to $448\times 448$ can achieve nontrivial performance boosts. However, further increasing the size to $896\times 896$ will harm the performances. This phenomenon indicates that excessively increasing the input size (to a scale dramatically different from its pretraining) will break the semantics of visual features, and thereby damaging the vision-language alignment.
C.2 Vision Encoders
@@
Since the input size can not be seamlessly scaled up as discussed above, it is worth further exploring to leverage vision encoders naturally adaptive to high-resolution image inputs. Toward this end, we employ SAM ViT-H/16 [34], which takes $1024\times 1024$ images as inputs, as an additional vision backbone to provide high-resolution visual features for V-L dense aligner. However, as shown in Table 9(b), the utilization of SAM ViT-H/16 results in a significant performance deterioration. We posit that this phenomenon can be attributed to the inadequate alignment between the SAM visual encoder and the language modality. Building upon the aforementioned experimental findings and analysis, we think that embracing vision encoders capable of processing high-resolution image inputs without compromising semantic coherence can pave the way for further enhancing the vision-centric capabilities of LMMs.
C.3 Number of K for Dense Prediction Tasks
We also investigate the impact of selecting different values for $K$ on the dense prediction tasks, e.g., object detection. As shown in Table 9(c), when $K$ increases from 20 to 100, the model’s performance consistently improves, which can be attributed to the increased recall of true positives. As $K$ further increases from 100 to 200, the model’s performance does not improve evidently, therefore, we set $K$ as 100 in our implementation.
C.4 Training Epochs
We record the variations in detection performance as the number of training iterations increases in Table 9(d). It can be observed that the model’s performance slowly increases as the training proceeds. Due to the limitation of computational resources, we did not train our model for a longer time. The convergence speed might be limited by the optimization difficulty of utilizing a single [LOC] token to query the entire image regions. A feasible solution to mitigate this problem is to employ more special tokens like [LOC] serving as object queries in DETR [52] with the Hungarian matching for label assignment, which can accelerate training and promote performances as proved in VisionLLM [36]. We do not adopt this design as it is unclear whether these tokens customized for object detection can generalize well to other tasks.
Table 9: Ablation studies on model designs.
(a) Effect of different input sizes for the dense alignment.
Input Size AP AP50 AP75
336$\times$336 26.1 40.2 25.6
448$\times$448 28.4 45.0 28.1
896$\times$896 24.5 36.4 22.5
(b) Effect of different pre-trained vision encoders.
@@ C.4 Training Epochs
SAM ViT-H/16-1024 14.5 24.0 14.1
CLIP ViT-L/14-336 28.4 45.0 28.1
(c) Effect of different $K$ value on dense prediction task.
@@
AP 26.3 27.6 28.4 28.5
AP50 39.5 43.2 45.0 45.3
AP75 25.3 27.2 28.1 28.0
(d) Effect of different numbers of training iterations.
@@
AP 28.4 31.6 34.5 35.2 35.3
AP50 45.0 49.2 52.4 53.2 53.2
AP75 28.1 31.5 34.7 35.5 35.8
Figure 6: Qualitative results of versatile vision-centric capabilities of our proposed Lumen. We use different colors of dotted rectangles to denote performing different tasks.
Appendix D More Qualitative Results
We provide qualitative results to demonstrate the versatile capabilities of our Lumen as shown in Fig 6. Overall, the abundant qualitative results prove that our method possesses versatile vision-centric capabilities. It is worth noting that in the case within the second row, we instruct the model to detect the non-existing object (i.e., “basketball” in Fig 6), the maximum activation across the heatmap is relatively low, and filtered with a pre-defined threshold. Therefore, the model will not generate any output for this wrong instruction as illustrated in Fig 6.